# LLM Fingerprinting via Semantically Conditioned Watermarks

### v2

@@ 1 Introduction
Training state-of-the-art Language Models (LMs) is an expensive process (Maslej et al., 2025). When releasing new open-weight LMs, companies (model owners) therefore adopt restrictive licenses, for example prohibiting commercial use. As third-party malicious deployers may deploy such a model behind an API without honoring the license, model owners need a reliable way to prove ownership.
@@ 2 Related Work
Model fingerprinting typically follows one of two approaches: white-box methods operating on weights/activations, and black-box methods relying only on model outputs, which are the focus of this work and the discussion in Section˜1. White-box fingerprints (Zhang et al., 2024, Zeng et al., 2023, Refael et al., 2024) have so far shown both robustness and stealthiness, however because detection requires access to the weights of the model, they are practically limited in most realistic scenarios. Black-box fingerprinting methods (Xu et al., 2024a, Zeng et al., 2024, Pasquini et al., 2024, Nasery et al., 2025) embed specific query-key backdoor into the model. We consider two prominent baselines from prior work: Instructional Fingerprinting (IF) (Xu et al., 2024a) and Scalable Fingerprinting (SF) (Nasery et al., 2025), detailing both methods in Section˜5.
@@ A Semantic Query Domain
To make it harder to block queries and disable the fingerprint by perturbing them (e.g., system prompts), we propose replacing the small finite set of queries with a broader semantic domain. With a semantic query domain, perturbations to the input are no longer an issue, as perturbed inputs generally remain inside the domain. In Section˜C.2, we find that it is necessary for the domain to have high entropy (i.e., the average entropy of the model distribution on this domain should be high) to ensure the text generated by the model contains the LLM watermark signal. At the same time, as discussed in Section˜5.3, the domain should be restrictive enough to prevent targeted adversaries from detecting the fingerprint. In this work, we use French as a domain for our main experiments (Section˜5) and explore additional domains in Sections˜C.2 and E.
@@ 4.1 Embedding Our Fingerprint
We present an overview of our fingerprint learning method in Algorithm˜1. At a high level, the model owner starts with the LLM $\theta$ to be fingerprinted, a semantic domain dataset $\mathcal{D}_{\text{target}}$, and the Red-Green watermark parameters, namely the context size $k$, the watermark strength $\delta$, and the private key $\xi$, introduced in Section˜2. The optimization objective then requires balancing the following goals: ensuring no distortion of the model distribution outside the semantic domain (regularization term $L_{\text{reg}}$, Line 6), and learning the watermark distribution on the semantic domain (watermark distillation term $L_{\text{watermark}}$, Line 5). We describe these two components in more detail shortly. In Sections˜5.1 and 5.2, we experimentally show that this method enables a robust signal with minimal impact on utility. We find in Section˜5.3 that this also ensures stealthiness against even adversarial and targeted adversaries and in Section˜B.3 that it also improves robustness against finetuning. We ablate the choice of the semantic domain in Section˜C.2 and the watermark hyperparameters in Section˜C.3.
@@ 4.2 Detecting Our Fingerprint
$Z(\omega)=\frac{\hat{\gamma}(\omega)-\gamma}{\sqrt{\beta(\omega)\gamma(1-\gamma)/|\omega|}},$ (3)
@@ 5.1 Fingerprint Effectiveness
Table˜1 shows that, for all three models, the fingerprint is systematically detected when the model is fingerprinted, and importantly, not detected for the base, non-fingerprinted model. Regarding utility, we find that across all $8$ benchmarks tested, the accuracies do not significantly drop. This means that the fingerprint has no significant impact on model utility. Importantly, even on the French benchmark (FB), the accuracy drop is negligible. These results validate our semantically conditioned watermarking approach, specifically our regularization loss in Eq.˜2 which preserves overall model performance. Indeed, in Section˜C.4, we find that embedding our fingerprint without regularization leads to a more significant drop in benchmark accuracies. In Section˜B.1, we use additional metrics, namely perplexity and LLM-as-a-judge scores (with GPT5-mini). We find that our fingerprint almost perfectly preserves the model distribution outside the semantic domain, and only distorts the distribution (due to the watermark) within the semantic domain. For completeness, we provide examples of model replies in App.˜G.
@@ 5.3 Fingerprint Stealth
Fig.˜2 shows the FPR and recall of the GPT5-mini judge classifier. With the FPR, we see that our judge is reliable: there are almost no false positives (<3<3%). We see that both queries and keys from IF are detected by the judge; hence, it is not stealthy. In contrast, for queries, both our approach and SF are completely undetected, and for the outputs, the recall of our approach and SF are relatively low. As both our fingerprint and SF use natural queries/keys, these results show that using natural text as fingerprints is essential for stealth. In our paradigm, by using a semantic domain, we ensure that the queries are stealthy. For the replies, while the watermark might distort the distribution of the text in detectable ways (e.g., watermarked LLMs can be detected via specific prompting (Gloaguen et al., 2025a, Liu et al., 2025)), these results show that (i) it is difficult for non-specific adversaries to detect the watermark and (ii) even for specific adversaries like Gloaguen et al. (2025a), they need to know the semantic domain beforehand to be effective. Thus, in practice, our fingerprint is stealthy.
@@ 6 Conclusion & Limitations
To ensure reproducibility, we clearly detail our fingerprinting embedding algorithm in Section˜4.1 and the corresponding fingerprint detection algorithm in Section˜4.2. For all experimental evaluations in Section˜5, we introduce the hyperparameters before each experiment and provide a comprehensive list in App.˜A. We also include our code with the submission.
@@ Appendix B Additional Experiments
In this section, we perform additional evaluation of our fingerprint. We evaluate its impact on model quality (Section˜B.1) using perplexity and GPT5-as-a-judge scores, and we evaluate its robustness against additional scenarios in Section˜B.2. In Section˜B.3, we show that, in addition to being stealthier, semantically conditioned watermarks are also more robust to finetuning than watermarking all model outputs.
@@ B.2 Additional Fingerprint Robustness Evaluation
We evaluate our robustness to system prompts on two leaked system prompts from GPT5. The Robot one corresponds to the Robot personality of GPT5 and the second one to the base personality. We used the system prompts as is but removed the parts related to tooling. Table˜5 shows that, with realistic system prompts used in production, only our fingerprint is completely robust and maintains a FSR of $1.0$.
@@ Appendix C Ablation on Our Fingerprinting Method Components
We ablate the different component of our method (Section˜4), namely the number of queries $|Q|$ required for detection (Section˜C.1), the watermark strength parameter $\delta$ used when embedding the fingerprint (Section˜C.3), the semantic domain on which the fingerprint is embedded (Section˜C.2) and the effect of the regularization loss from Eq.˜2 (Section˜C.4).
@@ Results
Fig.˜5 shows that detectability increases with the number of queries. For all three settings evaluated, we see that the fingerprint is detected with $100$ queries already. This highlights the strength of using a statistical signal: even in cases where the signal is degraded, we can recover a strong fingerprint signal simply by scaling the number of queries. Moreover, the base model fingerprint detectability (orange line) is flat, showing that our fingerprint is principled and does not induce an abnormal false positive rate.
@@ C.2 Ablation on the Semantic Domain
Table˜6 shows that for all semantic domains we achieve a Fingerprint Success Rate of $1.0$. Moreover, independently of the domain, we observe no significant drop in benchmark accuracies. These results indicate that our fingerprint has a minor impact on the model utility, even in more technical domains such as mathematics.
@@
Fig.˜6 shows that the detectability with respect to the number of queries scales differently across domains. This is because LLM watermark detectability (i.e., the detection power of the watermark) is bounded by the entropy of the LLM distribution (Kirchenbauer et al., 2023). Hence, the entropy of the underlying semantic domain is critical when choosing the domain. On the positive side, we see that for all evaluated domains increasing the number of queries indeed increases the fingerprint detectability.
@@ C.3 Ablation on the Watermark Strength Parameter
We embed our fingerprint in Qwen2.5-3B on the French domain using the same hyperparameters as described in Section˜5, but with a different watermark strength parameter $\delta$, ranging from $0.5$ to $4$. To evaluate fingerprint detectability, we query the model up to $1000$ times with prompts from the French Alpaca dataset. To evaluate the impact of $\delta$ on quality in the semantic domain, we compute the perplexity of the generated replies using Qwen2.5-32B.
@@ Results
Fig.˜7 shows that increasing $\delta$ indeed significantly increases detectability, and that $1000$ queries are enough to detect the fingerprint for all $\delta$ tested. The right side of Fig.˜7 shows that, at the same time, increasing $\delta$ also slightly increases perplexity in the semantic domain. This means that, for the model provider, there is a trade-off between the cost of detectability (due to the increased number of queries needed to detect the watermark) and the impact on quality. Given that the impact on quality is minimal even with $\delta=4$ (see the benchmark results from Section˜5.1) and limited to the semantic domain, we argue that detectability is more important and thus decided to use $\delta=4$ for our main experiments in Section˜5
@@ C.4 Ablation on Our Fingerprint Regularization Loss
Table˜7 shows the effectiveness of our fingerprint with and without regularization. While the absence of regularization does not alter the detectability of the fingerprint (FSR remains at 1), we notice a slight degradation in model utility despite the watermark distillation loss being applied only on the French dataset. This already highlights the importance of the regularization term in our fingerprinting method. Additionally, as mentioned in Section˜5.3, using a semantic domain is important for stealth as it prevents targeted watermark detection attacks (Gloaguen et al., 2025a, Liu et al., 2025).

### v3

@@ 1 Introduction
Training state-of-the-art Language Models (LMs) is an expensive process (Maslej et al., 2025). When releasing new open-weight LMs, companies (model owners) therefore adopt restrictive licenses, for example prohibiting commercial use. As third-party malicious deployers may deploy such a model behind an API without honoring the license, model owners need a reliable way to prove ownership. Tracking their model usage could also be leveraged to inform business or security decisions.
@@ 2 Related Work
Model fingerprinting typically follows one of two approaches: white-box methods operating on weights/activations, and black-box methods relying only on model outputs, which are the focus of this work and the discussion in Section˜1. White-box fingerprints (Zhang et al., 2024, Zeng et al., 2023, Refael et al., 2024) have so far shown both robustness and stealthiness, however because detection requires access to the weights of the model, they are practically limited in most realistic scenarios. Black-box fingerprinting methods (Xu et al., 2024a, Zeng et al., 2024, Pasquini et al., 2024, Nasery et al., 2025) embed specific query-key backdoor into the model. We consider two prominent baselines from prior work: Instructional Fingerprinting (IF) (Xu et al., 2024a) and Scalable Fingerprinting (SF) (Nasery et al., 2025), detailing both methods in Section˜5. The main motivation for model fingerprinting is to give model providers a tool to enforce their licenses. While no precedents have been established regarding whether such licenses are enforceable, model ownership protection (Anthropic, 2025) and open-source LLM licenses play important roles for companies. Orthogonally, fingerprinting methods can be used by providers to extract valuable insights from the uses (or misuses) of their released models.
Alternatives to Fingerprinting
Another line of work comprises timing attacks on LLM APIs (Carlini & Nasr, 2024, Song et al., 2025, Alhazbi et al., 2025), where several properties of the deployed model are inferred directly from the API response time. Although such techniques could be used as a fingerprinting method to uniquely identify a model, Carlini & Nasr (2024) in particular show that strong defenses against such attacks exist. We believe this is an interesting direction for future work.
@@ A Semantic Query Domain
To make it harder to block queries and disable the fingerprint by perturbing them (e.g., system prompts), we propose replacing the small finite set of queries with a broader semantic domain. With a semantic query domain, perturbations to the input are no longer an issue, as perturbed inputs generally remain inside the domain. In Section˜C.2, we find that it helps for the domain to have high entropy (i.e., the average entropy of the model distribution on this domain should be high) to ensure the text generated by the model contains the LLM watermark signal. Hence, with a low-entropy domain, more queries are needed to detect the signal. At the same time, as discussed in Section˜5.3, the domain should be restrictive enough to prevent targeted adversaries from detecting the fingerprint. In this work, we use French as a domain for our main experiments (Section˜5) and explore additional domains in Sections˜C.2 and E.
@@ 4.1 Embedding Our Fingerprint
We present an overview of our fingerprint learning method in Algorithm˜1. At a high level, the model owner starts with the LLM $\theta$ to be fingerprinted, a semantic domain dataset $\mathcal{D}_{\text{target}}$, and the Red-Green watermark parameters, namely the context size $k$, the watermark strength $\delta$, and the private key $\xi$, introduced in Section˜2. The optimization objective then requires balancing the following goals: ensuring no distortion of the model distribution outside the semantic domain (regularization term $L_{\text{reg}}$, Line 6), and learning the watermark distribution on the semantic domain (watermark distillation term $L_{\text{watermark}}$, Line 5). We describe these two components in more detail shortly. In Sections˜5.1 and 5.2, we experimentally show that this method enables a robust signal with minimal impact on utility. We find in Section˜5.3 that this also ensures stealthiness against even adversarial and targeted adversaries and in Section˜B.5 that it also improves robustness against finetuning. We ablate the choice of the semantic domain in Section˜C.2 and the watermark hyperparameters in Section˜C.3.
@@ 4.2 Detecting Our Fingerprint
$Z(\omega)=\frac{\hat{\gamma}(\omega)-\gamma}{\beta(\omega)\sqrt{\gamma(1-\gamma)/|\omega|}},$ (3)
@@ 5.1 Fingerprint Effectiveness
Table˜1 shows that, for all three models, the fingerprint is systematically detected when the model is fingerprinted, and importantly, not detected for the base, non-fingerprinted model. Regarding utility, we find that across all $8$ benchmarks tested, the accuracies do not significantly drop, except for Llama3.2-1B on HumanEval. Given that smaller models tend to be overfinetuned and sensitive to perturbations of the weights (Springer et al., 2025), we believe that this drop on HumanEval is independent of our method and is instead due to the overall stronger impact of finetuning smaller models. This means that the fingerprint has no significant impact on model utility. Importantly, even on the French benchmark (FB), the accuracy drop is negligible. These results validate our semantically conditioned watermarking approach, specifically our regularization loss in Eq.˜2 which preserves overall model performance. Indeed, in Section˜C.4, we find that embedding our fingerprint without regularization leads to a more significant drop in benchmark accuracies. In Section˜B.1, we use additional metrics, namely perplexity and LLM-as-a-judge scores (with GPT5-mini). We find that our fingerprint almost perfectly preserves the model distribution outside the semantic domain, and only distorts the distribution (due to the watermark) within the semantic domain. For completeness, we provide examples of model replies in App.˜G.
@@ 5.3 Fingerprint Stealth
Fig.˜2 shows the FPR and recall of the GPT5-mini judge classifier. With the FPR, we see that our judge is reliable: there are almost no false positives (<3<3%). We see that both queries and keys from IF are detected by the judge; hence, it is not stealthy. In contrast, for queries, both our approach and SF are completely undetected, and for the outputs, the recall of our approach and SF are relatively low. As both our fingerprint and SF use natural queries/keys, these results show that using natural text as fingerprints is essential for stealth. In our paradigm, by using a semantic domain, we ensure that the queries are stealthy. For the replies, while the watermark might distort the distribution of the text in detectable ways (e.g., watermarked LLMs can be detected via specific prompting (Gloaguen et al., 2025a, Liu et al., 2025)), these results show that (i) it is difficult for non-specific adversaries to detect the watermark and (ii) even for specific adversaries like Gloaguen et al. (2025a), they need to know the semantic domain beforehand to be effective. We verify in Section˜C.5 that the fingerprint can indeed only be detected on the semantic domain. Thus, in practice, our fingerprint is stealthy.
@@ 6 Conclusion & Limitations
To ensure reproducibility, we clearly detail our fingerprinting embedding algorithm in Section˜4.1 and the corresponding fingerprint detection algorithm in Section˜4.2. For all experimental evaluations in Section˜5, we introduce the hyperparameters before each experiment and provide a comprehensive list in App.˜A. Our code is available here.
@@ Appendix B Additional Experiments
In this section, we perform additional evaluation of our fingerprint. We evaluate its impact on model quality (Section˜B.1) using perplexity and GPT5-as-a-judge scores, and we evaluate its robustness against additional scenarios in Section˜B.2. In Section˜B.5, we show that, in addition to being stealthier, semantically conditioned watermarks are also more robust to finetuning than watermarking all model outputs.
@@ B.2 Additional Fingerprint Robustness Evaluation
Pirate 0.0 0.0 1.0 0.0 0.0 1.0 0.0 0.8 1.0
Weather 0.0 0.0 1.0 0.0 0.0 1.0 0.0 1.0 1.0
Active Input Translation* 0.0 0.0 1.0 0.0 0.0 1.0 0.0 0.0 1.0
Output Translation* 0.0 0.4 1.0 0.0 0.0 1.0 0.0 0.0 1.0
@@
We evaluate our robustness to system prompts on two leaked system prompts from GPT5 and two specially designed adversarial system prompts. The Robot one corresponds to the Robot personality of GPT5 and the second one to the base personality. We used the system prompts as is but removed the parts related to tooling. The two additional prompts, Pirate and Weather, are listed below.
  1. 1.
    Pirate: Talk like a pirate.
  2. 2.
    Weather: Only answer weather-related queries.
Table˜5 shows that, with realistic system prompts used in production and adversarial system prompts, only our fingerprint is completely robust and maintains a FSR of $1.0$. For adversarial system prompts in particular, we see that the baselines almost systematically fail, whereas our approach consistently reaches an FSR of 1.0.
@@
Advanced Back-translation
We further evaluate the robustness of our fingerprint against active adversaries. Specifically, for both input and output, we consider an advanced back-translation attacker (Translation*) that uses two different models (compared to only one in Section˜5.2). In particular, we translate from the prompt language to Chinese using GPT5-mini and then from Chinese back to the prompt language using Gemini2.5-Flash-Lite. We find in Table˜5 that our fingerprint consistently reaches an FSR of $1.0$ against such an adversary, unlike both baselines.
@@ B.3 Fingerprinting With Full Watermark
B.4 Robustness to Instruction-Tuning
In this section, we evaluate whether (i) we can use our fingerprint on base models and (ii) whether our fingerprint is robust to instruction-tuning (and a subsequent change in prompt format).
Figure 5: Robustness to Instruction-Tuning We show the detectability of our fingerprint (measured by the negative log p-value) with respect to the number of queries $|Q|$ after instruction-tuning. We generate the replies with Qwen2.5-3B, and average all results over $5$ independent runs. In red, we show the fingerprint decision threshold of 11e03-03.
Setup
For embedding our fingerprint (Section˜4.1) in a base model, we use the same hyperparameters as those described in App.˜A, but change the semantic domain dataset $\mathcal{D}_{\text{target}}$ and the regularization dataset $\mathcal{D}_{\text{reg}}$ to completion ones. In particular, we embed our fingerprint in the Health domain and use the health subcategory of the WebOrganizer dataset (Wettig et al., 2025) as the target dataset. For the regularization dataset, we also use the WebOrganizer dataset but without the health subcategory. As the model, we use the completion version of Qwen2.5-3B instead of the instruction-tuned one. Then, we instruction-tuned our fingerprinted model using the Qwen chat template on Alpaca. Specifically, we finetune for one epoch with a batch size of $64$, a learning rate of 22e05-05, a warmup of 1010% with a cosine learning rate scheduler, and the Adafactor optimizer. To evaluate the robustness of our fingerprint, we then use $1000$ health-related questions extracted from the report in WHO, Regional Office for Africa (2019), using the newly learnt chat template.
Robustness Against Instruction-Tuning
Fig.˜5 shows that our fingerprint is robust to instruction-tuning and effective on completion models. Indeed, with $1000$ queries, the FSR is consistently 1.0. This highlights the inherent robustness of both using a semantic domain as fingerprint queries and a statistical signal as keys.
B.5 Evaluation of Our Fingerprint Impact on Finetuneability
In this part, we measure the impact of our fingerprint on finetuneability, i.e., whether it is harder to improve performance through finetuning a fingerprinted model compared to the original version.
Table 6: Evaluation of Our Fingerprint Impact on Finetuneability We compare the benchmark performance of Qwen2.5-3B models (without/with our fingerprint on French) on task-specific benchmarks (respectively GalicianBench and FrenchBench) before and after finetuning on the corresponding datasets (respectively AlpacaGalician and WildChatFr).
Model GalicianBench FrenchBench
Before After Before After
Qwen2.5-3B 46 50 59 63
Qwen2.5-3B (Fingerprinted) 46 51 57 59
Setup
We compare our Qwen2.5-3B fingerprinted model on the French domain with the original version. We independently evaluate finetuneability within the semantic domain and outside the semantic domain. Within the semantic domain, we finetune both models on WildChatFr for one epoch (using a learning rate of 22e5-5, a batch size of $64$, and the Adafactor optimizer with a cosine scheduler and a warmup ratio of 1010%). We then measure the benchmark accuracy on FrenchBench before and after finetuning. Outside the semantic domain, we follow a similar procedure by finetuning both models on AlpacaGalician (Bao et al., 2023) (using the same hyperparameters but for $3$ epochs). We then measure the benchmark accuracy on GalicianBench (Baucells et al., 2025) (GB).
Results
Table˜6 shows that both without and with our fingerprint, finetuning indeed improves the model performance on the corresponding task. When finetuning outside the semantic domain (i.e., on Galician), we observe that the performance before and after finetuning of the fingerprinted model is the same as that of the original model. Within the semantic domain (i.e., on French), although the performance of the fingerprinted model is lower, it also increases. Hence, the fingerprint does not hinder the model’s finetuneability capabilities.
@@ Appendix C Ablation on Our Fingerprinting Method Components
We ablate the different component of our method (Section˜4), namely the number of queries $|Q|$ required for detection (Section˜C.1), the watermark strength parameter $\delta$ used when embedding the fingerprint (Section˜C.3), the semantic domain on which the fingerprint is embedded (Section˜C.2), the effect of the regularization loss from Eq.˜2 (Section˜C.4 and Section˜C.5).
@@ Results
Fig.˜6 shows that detectability increases with the number of queries. For all three settings evaluated, we see that the fingerprint is detected with $100$ queries already. This highlights the strength of using a statistical signal: even in cases where the signal is degraded, we can recover a strong fingerprint signal simply by scaling the number of queries. Moreover, the base model fingerprint detectability (orange line) is flat, showing that our fingerprint is principled and does not induce an abnormal false positive rate.
@@ C.2 Ablation on the Semantic Domain
Table˜7 shows that for all semantic domains we achieve a Fingerprint Success Rate of $1.0$. Moreover, independently of the domain, we observe no significant drop in benchmark accuracies. These results indicate that our fingerprint has a minor impact on the model utility, even in more technical domains such as mathematics.
@@
Fig.˜7 shows that the detectability with respect to the number of queries scales differently across domains. This is because LLM watermark detectability (i.e., the detection power of the watermark) is bounded by the entropy of the LLM distribution (Kirchenbauer et al., 2023). Hence, the entropy of the underlying semantic domain is critical when choosing the domain. On the positive side, we see that for all evaluated domains increasing the number of queries indeed increases the fingerprint detectability.
@@ C.3 Ablation on the Watermark Strength Parameter
We embed our fingerprint in Qwen2.5-3B on the French domain using the same hyperparameters as described in Section˜5, but with a different watermark strength parameter $\delta$, ranging from $0.5$ to $4$. To evaluate fingerprint detectability, we query the model up to $1000$ times with prompts from the French Alpaca dataset. To evaluate the impact of $\delta$ on quality in the semantic domain, we compute the perplexity of the generated replies using Qwen2.5-32B, and the benchmark accuracy on FrenchBench.
@@ Results
Fig.˜8 shows that increasing $\delta$ indeed significantly increases detectability, and that $1000$ queries are enough to detect the fingerprint for all $\delta$ tested. The right side of Fig.˜8 shows that, at the same time, increasing $\delta$ also slightly increases perplexity in the semantic domain. Similarly, increasing $\delta$ also slightly decreases FrenchBench accuracy (from $0.61$ at $\delta=0.5$ to $0.57$ at $\delta=5$). This means that, for the model provider, there is a trade-off between the cost of detectability (due to the increased number of queries needed to detect the watermark) and the impact on quality. Given that the impact on quality is minimal even with $\delta=4$ (see the benchmark results from Section˜5.1) and limited to the semantic domain, we argue that detectability is more important and thus decided to use $\delta=4$ for our main experiments in Section˜5
@@ C.4 Ablation on Our Fingerprint Regularization Loss
Table˜8 shows the effectiveness of our fingerprint with and without regularization. While the absence of regularization does not alter the detectability of the fingerprint (FSR remains at 1), we notice a slight degradation in model utility despite the watermark distillation loss being applied only on the French dataset. This already highlights the importance of the regularization term in our fingerprinting method. Additionally, as mentioned in Section˜5.3, using a semantic domain is important for stealth as it prevents targeted watermark detection attacks (Gloaguen et al., 2025a, Liu et al., 2025).
C.5 Ablation on Fingerprint Leakage
In this part, we show that the fingerprint can only be detected in the semantic domain, and that querying the fingerprinted model on other domains does not leak the fingerprint. This is important to ensure the stealth benefits of semantically conditioning the fingerprint, detailed in Section˜5.3.
Figure 9: Ablation on Fingerprint Leakage We compare the detectability of our fingerprint (measured by the negative log p-value) with respect to the number of queries $|Q|$ for different semantic domains. We generate the replies with Qwen2.5-3B fingerprinted on the math domain, and average all results over $5$ independent runs. In red, we show the fingerprint decision threshold of 11e03-03. We see that only with the math domain is the fingerprint detected: there is no leakage.
Setup
We use a similar setup as in Section˜C.2. We embed on Qwen2.5-3B our fingerprint using math as a semantic domain but without the regularization loss (Eq.˜2). For detection, we use $|Q|=1000$ different queries from the GSM8K benchmark (Math). To test for fingerprint leakage, we also use $|Q|=1000$ different queries from the CodeAlpaca (Chaudhary, 2023) dataset (Code), $|Q|=1000$ different health-related queries extracted from WHO, Regional Office for Africa (2019) (Medicine), and $|Q|=1000$ different queries from the Alpaca (Taori et al., 2023) dataset (Q&A).
No Fingerprint Leakage
Fig.˜9 shows that there is no fingerprint leakage. Even when using $1000$ queries for fingerprint detection, only the queries in the semantic domain, i.e., math, contain a detectable watermark signal. This means that, for a targeted adversary to detect the fingerprint (such as in Gloaguen et al. (2025a)), they would indeed need to know the semantic domain beforehand.